if $100 is put in a bank account with a 5% interest rate, the money will become

Pt=100⋅1+0.05t

after t years. How long will it take for the money to double?
Answer #1

Solutions Given that britially P $100 Plt) = 100/1+0.5+) r=57. for. Double money. 200 = 100+ o(1+0.5t) 2= 1+0.5t i = 0.5t t=